Exercise 11-11

The stockholders’ equity section of Monty Corp. at December 31 is as follow

Monty Corp.
Balance Sheet(Partial)

Paid-in capital
   Preferred stock, cumulative, 14,000 shares authorized, 9,800 shares issued and outstanding

$588,000

   Common stock, no par, 880,000 shares authorized, 560,000 shares issued

1,120,000

     Total paid-in capital

1,708,000

Retained earnings

1,858,000

        Total paid-in capital and retained earnings

3,566,000

Less: Treasury stock (12,000 common shares)

62,400

             Total stockholders’ equity

$3,503,600

From a review of the stockholders’ equity section, as chief accountant, write a memo to the president of the company answering the following questions.

(a) How many shares of common stock are outstanding?

shares
(b) Assuming there is a stated value, what is the stated value of the common stock? $

per share
(c) What is the par value of the preferred stock? $

per share
(d) If the annual dividend on preferred stock is $52,920, what is the dividend rate on preferred stock?

%
(e) If dividends of $105,840 were in arrears on preferred stock, what would be the balance in Retained Earnings?
Answer #1

a)

Number of common shares outstanding = Number of common shares issued - Number of treasury shares

= 560,000 - 12,000

= 548,000

b)

Stated value of the common stock = Common stock paid in capital/ Number of common shares issued

= 1,120,000/560,000

= $2 per share

c)

Par value of the preferred stock = Preferred stock paid in capital/Number of preferred shares issued

= 588,000/9,800

= $60

d)

Dividend per preferred stock =  Annual dividend on preferred stock/Number of preferred shares issued

= 52,920/9,800

= $5.4

Dividend rate on preferred stock = Dividend per preferred stock/Par value of the preferred stock

= 5.4/60

= 9%

e)

Retained earnings balance = Retained earnings - Arrears of preferred dividends

= 1,858,000 - 105,840

= $1,752,160

Please ask if you have any query related to the question. Thank you